∫(e x /1+e 2x ) dx

∫(ex/1+e2x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{e^x}{e^{2x}+1}dx
I = \int \frac{e^x}{(e^{x})^2+1}dx
e^x = t
e^xdx = dt
I = \int \frac{1}{t^2+1}dt
I = tan^{-1}t+c
I = tan^{-1}(e^x)+c